A Mixture Contains 1.5 Liters Oxalic Acid H2C2O4· 2H2O (ρ=1.653 g/mL) and 2-

gallon Water. Determine:

ρ Oxalic acid = 1.653 g/mL

ρ water = 1000 g/L


a) %w/w, %w/v, %v/v of oxalic acid in a solution (6 pts)

b) ppm and ppb of oxalic acid (4 pts)

c) Molarity, Normality, and molality of the solution (10 pts)


1
Expert's answer
2021-10-20T02:50:57-0400

Molecular weight of oxalic acid (H2​C2​O4.2H2O) is 126 g/mol.


Molarity of 0.2 N oxalic acid (H2​C2O4.2H2O) solution "=\\frac{0.2\u200b}{2}=0.1 M"


Amount of solution =50 mL =0.05 L


So, amount of oxalic acid =126×0.1×0.05=0.63 g